# Design of a LYSO Crystal Electromagnetic Calorimeter for DarkSHINE Experiment

### v1

@@ III.2 Signal and background
In summary, the ECAL effectively excludes all soft backgrounds and hard backgrounds that do not involve muons or hadrons.
III.3 Segmentation optimization
Figure 8: (a) Structure of the ECAL viewed from the side. (b) Detailed structure of the ECAL for cell optimization, showing the placement of LYSO crystals (light blue), SiPMs (red), and the supporting PCB (green) and copper (gold) sheets used for electronic signal readout and cooling.
To thoroughly understand how crystal size influences energy measurement, a 1.2 mm thick PCB and a 0.5 mm thick copper plate were integrated behind each crystal layer for electronic signal readout and cooling purposes. Figure 8 depicts the detailed structure of the ECAL. Crystals are arranged in multiple layers along the beam direction, each layer consisting of uniformly positioned crystals. Each crystal is wrapped with an 80 $\mu$m ESR film and coupled with a 0.6 mm thick, 3×3 mm2 silicon piece at the end. A 20 $\mu$m air gap separates all materials. The ECAL comprises several thousand to tens of thousands of crystals, with the size affecting various performance aspects, such as light collection efficiency and the amount of passive material introduced. This study optimizes the detector’s performance by adjusting the crystal size to achieve high energy resolution and efficient energy containment.
Figure 9: (a) Light yield measured by SiPM from Geant4 optical simulation. (b) Maximum deposited energy within crystals under an 8 GeV electron beam incident on the ECAL. Crystals with smaller cross-sectional areas and shorter lengths exhibit higher light yields due to greater light collection efficiency, but this also results in lower energy absorption.
Figure 9 shows the simulated light yield and maximum deposited energy in LYSO crystals of various sizes. Crystals with smaller cross-sectional areas and shorter lengths exhibit higher light yields but lower deposited energy under an 8 GeV electron beam. The deposited energy was obtained from the full ECAL simulation. However, ECALs equipped with small crystals may contain more passive material due to the increased number of layers. Light yield and deposited energy influence the light output and the dynamic range of the electronics, while the amount of passive material impacts energy containment. These effects are quantified using the digitization tool.
Figure 10: (a) Equivalent noise energy (ENE) of a single channel for different setups. Larger crystals exhibit higher ENE, primarily due to their lower light yield. Additionally, larger crystals have a broader dynamic range for energy deposition. (b) Energy resolution of a single channel for different setups, incorporating the three digitization processes. Units with smaller crystals achieve better energy resolution.
Figure 10 illustrates the Equivalent Noise Energy (ENE) for crystals of different sizes. The ENE is calculated using the formula:
$ENE=\frac{Noise}{ADC_{p.e.}\times LY_{mea}}$ (1)
where $Noise$ is the number of ADC channels corresponding to electronic noise. For crystals of different sizes, the $Noise$ was set to a constant value of three ADC. $ADC_{p.e.}$ is the number of ADC channels per photoelectron, which varies with gain. $LY_{mea}$ is the light yield measured by SiPM, similar with the values in Figure 9(a). Parameters of $Noise$ and $ADC_{p.e.}$ for one specific setup have already been shown in Table 2, while $LY_{mea}$ is equivalent to the parameter $Lightyield(measured)$ in Table 1. Thus, during the digitization process, different crystal sizes have varying parameters.
As a result, in Figure 10(a), crystals of different sizes exhibit varying ENE values due to differences in light yields. Crystals with higher light yields have smaller ENE. For crystals of the same size, there are three distinct regions in the ENE corresponding to different gain modes. At the transition points between these ranges, the ENE exhibits jumps. Additionally, the dynamic range of energy deposition varies for different configurations. To ensure a balance between maintaining measurement integrity and achieving high measurement precision, we set different ADC dynamic ranges for each configuration, according to the product of the light yield and the maximum deposit energy. After incorporating the three digitization effects in  III.1.2, the single-channel energy resolution for these configurations is shown in Figure 10(b). Generally, the single-channel energy resolution improves as the crystal size decreases. The transition point around 500 MeV is due to the ENE jump caused by the range switching.
As a result, in Figure 10(a), crystals of different sizes exhibit varying ENE values due to differences in light yields. Crystals with higher light yields have smaller ENE. For crystals of the same size, there are three distinct regions in the ENE corresponding to different gain modes. At the transition points between these ranges, the ENE exhibits jumps. Additionally, the dynamic range of energy deposition varies for different configurations. To ensure a balance between maintaining measurement integrity and achieving high measurement precision, different ADC dynamic ranges were set for each configuration, based on the product of the light yield and the maximum deposited energy. After incorporating the three digitization effects in  III.1.2, the single-channel energy resolution for these configurations is shown in Figure 10(b). Generally, the single-channel energy resolution improves as the crystal size decreases. The transition point around 500 MeV is due to the ENE jump caused by the range switching.
Figure 11: Energy containment and energy resolution of the ECAL equipped with crystals of different sizes after digitization: (a) Ratio of deposited energy in the ECAL to the truth energy. Smaller crystals in the ECAL introduce a greater amount of passive material, resulting in lower energy containment. (b) Energy resolution, showing that ECALs equipped with small crystals have a better energy resolution below 4 GeV.
The performance of the full detector can be determined by applying digitization to each channel. Figure 11(a) and (b) show the energy containment and energy resolution of the ECAL equipped with crystals of different sizes. In these simulations, the overall size of the ECAL was kept around 52$\times$52$\times$48 cm3. The incident particles are electrons with energies ranging from 1 to 8 GeV, hitting the ECAL’s front face within a circular area with a radius of 3 cm at the center. Figure 11(a) shows that if the ECAL is composed of long and thick crystals, its energy containment improves. However, crystal length is more significant than cross-sectional area, as passive material is mainly concentrated in the PCB and copper layers between the layers. For energy resolution, when the ECAL is equipped with crystals of the same cross-sectional area (2.5$\times$2.5 cm2) but different lengths, shorter crystals result in better energy resolution. Conversely, when the ECAL is equipped with crystals of the same length (4 cm) but different cross-sectional areas, smaller cross-sectional areas lead to better energy resolution.
Because dark photon signals are often accompanied by significant energy loss in the ECAL, we are more concerned with the low-energy region below 4 GeV. Overall, when the ECAL is equipped with 2.5$\times$2.5$\times$4 cm3 crystals, it achieves high energy resolution for incident electron energies below 4 GeV while maintaining relatively balanced energy absorption efficiency. Additionally, crystals of this size provide good segmentation, offering potential for multivariate analysis, while keeping the number of channels manageable. Therefore, the crystal size of 2.5$\times$2.5$\times$4 cm3 is optimal for the DarkSHINE ECAL.
III.4 Volume optimization
@@
As the mass of dark photons increases, they carry more energy from the incident electron, resulting in a larger recoil angle. Figure 12 illustrates the distribution of recoil angles for particles striking the front surface of the ECAL for both inclusive events and dark photon processes with varying masses. The ECAL’s cross-sectional area is 52.5$\times$52.5 cm2. The majority of inclusive background processes involve electrons directly traversing the target and striking the central region of the ECAL, despite some divergence in the beam spot and slight displacement of electrons due to the magnetic field. In contrast, signal processes exhibit more dispersed recoil angles. Moreover, with increasing dark photon mass, more events occur in the peripheral region, indicating a higher likelihood of missing the ECAL. These events can be captured and vetoed by the HCAL. Therefore, increasing the ECAL volume can improve signal efficiency, though cost considerations must also be addressed.
@@
Figure 12: Dark photon signals exhibit larger recoil angles compared to inclusive background processes. As the mass of the dark photon increases, more events tend to hit on the peripheral regions of the ECAL.
@@
A signal region for the calorimeter is defined as shown in Table 4: the total energy in the ECAL is less than 2.5 GeV, the total energy in the HCAL is less than 30 MeV, and the maximum energy of a single unit in the HCAL is less than 0.1 MeV. This signal region is designed to achieve low background within $3\times 10^{14}$ EOTs. It is derived from the combined analysis of all sub-detectors, using thresholds for the ECAL and other sub-detectors, and applying extrapolation methods to exclude all backgroundsDarkSHINE ; wang2024design . The optimization results of the ECAL are evaluated by the number of events entering the signal region.
Table 4: Signal box for calorimeters.
@@
Figure 13: Signal efficiency as a function of dark photon mass for various ECAL configurations. (a) Optimization of the transverse size of the ECAL, showing signal efficiency for different transverse dimensions. It demonstrates that increasing the ECAL’s transverse size enhances signal efficiency, particularly for higher dark photon masses. (b) Optimization of the longitudinal size of the ECAL, showing signal efficiency for different longitudinal dimensions. It shows that increasing the ECAL’s longitudinal size improves signal efficiency.
To maintain consistency with the prospective studyDarkSHINE , we only use the truth energy here instead of the digitized energy. Additionally, the PCB and copper layers are not included in the geometry here and in subsequent analyses.
The ECAL size was optimized in two dimensions: transverse and longitudinal. The individual crystal size remains constant at 2.5$\times$2.5$\times$4.0 cm3, as optimized in section III.3. First, the ECAL’s transverse dimensions were adjusted while keeping the longitudinal length constant at 11 layers. Four different calorimeter sizes were simulated. Results in Figure 13(a) show that as the transverse size increases, signal efficiency for different dark photon masses also increases. The average signal efficiency for all mass points under each size is listed in Table 5.
@@
IV Energy distribution
As a fixed target experiment, incident particles without significant energy loss primarily hit the central area of the ECAL, resulting in substantial energy deposition in the central region’s crystals in the first few layers. Additionally, the energy absorbed by crystals in different regions varies greatly, with those closer to the edges typically experiencing minimal energy deposition. To determine the energy dynamic range of the future detector and explore potential triggering methods for the ECAL, the energy deposition in crystals from different regions was investigated.
IV.1 Energy deposition in different regions
With the optimized ECAL configuration in III.3 and III.4, the energy deposition in crystals for dark photon signals and background processes is shown in Figure 14. Inclusive background processes result in larger energy deposition in the crystals, up to 4 GeV. Energy deposition for dark photon signals is slightly lower, with smaller mass dark photons depositing more energy. The dark photon process with a 1 MeV dark photon mass can deposit up to 3.5 GeV of energy in the crystals, as dark photons have a very low probability of carrying minimal kinetic energy, resulting in the recoiling electron retaining most of its energy.
Figure 14: Energy deposited in single crystals under 8 GeV electron beam. This figure indicates that for inclusive background processes, the energy absorbed by a single crystal can reach up to 4 GeV, while for dark photon signal processes, it can reach up to 3.5 GeV.
Figure 15: ECAL divided into four regions, from the center of the shower to the periphery. Red: region-I, green: region-II, blue: region-III, yellow: region-IV.
@@ IV.1 Energy deposition in different regions
The ECAL was divided into four sections (Figure 15) from the center of the shower to the periphery to study energy deposition in different regions. The number of crystals in each region is shown in Table 7. Region-I (red) is the core area of shower development, containing 125 crystals each with a volume of 2.5$\times$2.5$\times$4 cm3, approximately equivalent to 18 radiation lengths. It covers the entire beam spot with a radius of 3 cm and an additional area of one Molière radius. The energy absorbed by each crystal in these four regions is shown in Figure 16. Crystals in the central and near-central regions absorb significantly more energy. In contrast, crystals in Region-III and Region-IV, farther from the shower center, absorb noticeably less energy. For dark photon signals, the energy deposition is even lower than for inclusive background. Figure 17 shows that crystals in Region-I, with only 2.6% of the volume, absorb more than 90% of the energy for inclusive backgrounds.
@@
Figure 16: Energy absorbed by each crystal in the four regions shown in Figure 15. Crystals in the central regions (Region-I and Region-II) absorb significantly more energy, while the crystals in Region-III and Region-IV, farther from the shower center, absorb noticeably less energy.
Figure 17: Energy ratio in different regions for different processes. Both signal and background processes deposit more energy in Region-I. However, for dark photon signal processes, as the dark photon mass increases, more energy is absorbed in the outer regions.
IV.2 Energy limits on crystal
The dynamic range setting affects the accuracy of energy measurement. Generally, a smaller energy dynamic range enables higher measurement accuracy. To determine the minimum dynamic range that meets the energy measurement requirements of the ECAL, an energy limit was applied to each crystal. When the energy deposited in the crystal exceeds these limit values, it is capped at the limit value, simulating the saturation behavior of a real detector. Energy limitations might reduce the total energy measured by the ECAL. Since interest lies in events with significant energy loss, the dynamic range is selected based on the number of events where the total energy measured by the ECAL is less than 4 GeV after applying the energy limits.
Table 8: Ratio of events with total energy less than 4 GeV under different crystal energy limitations.
@@ IV.2 Energy limits on crystal
Figure 18: Energy distribution after applying energy limits on crystals. When the energy limit is too low, background events can enter the signal box. However, when the energy limit reaches 4 GeV, the total energy closely matches the true value.
Table 8 shows the ratio of events with total energy less than 4 GeV for the background and dark photon signals of different masses after applying various crystal energy limits. For one million background processes, when the energy limit is set to 1 GeV, almost no events pass the cut, except for two photon-muon pair processes and one nuclear process, which can be vetoed by the HCAL. This indicates that very few crystals have an energy deposition exceeding 1 GeV. However, as shown in Figure 18, the total energy measured by the ECAL is significantly lower than the true value at this limit. Additionally, as the number of events increases, there should be a higher probability of background events entering the sensitive range. From Figure 18, a more suitable energy limit appears to be 4 GeV. At this limit, the total energy measured by the ECAL closely matches the true value, and it is difficult for background cases to enter the sensitive range through saturation. Therefore, a single crystal energy dynamic range of 4 GeV is sufficient for the DarkSHINE ECAL. The readout electronics in future detectors can be designed based on this standard.
IV.3 Trigger efficiency
Since the majority of events collected by the ECAL are background processes that are not of interest, data transmission and storage pressure can be alleviated by implementing trigger strategies based on the energy distribution within the ECAL. Using the criterion of energy less than 4 GeV for event selection, if the sum of the energy in certain crystals exceeds 4 GeV in an event, the event will not be saved. Given that crystals in Region-I absorb most of the energy from the recoiling electron, this region should exhibit the highest trigger efficiency.
Table 9: Ratio of events with region energy less than 4 GeV.

### v2

@@ III.2 Signal and background
In summary, the ECAL aims at excluding backgrounds that do not involving hard muons or hadrons.
IV Volume optimization
@@
As the mass of dark photons increases, they carry more energy from the incident electron, resulting in a larger recoil angle, which is defined as the angle between the particle momentum direction and the beam direction. Figure 6 illustrates the distribution of recoil angles for particles striking the front surface of the ECAL for both inclusive events and dark photon processes with varying masses, where the ECAL’s cross-sectional area is 52.5$\times$52.5 cm2. The majority of inclusive background processes involve electrons directly traversing the target and striking the central region of the ECAL, despite some divergence in the beam spot and slight displacement of electrons due to the magnetic field. In contrast, signal processes exhibit more dispersed recoil angles. Moreover, with increasing dark photon mass, more events occur in the peripheral region, indicating a higher likelihood of missing the ECAL. But these events can be captured and vetoed by the HCAL. Therefore, increasing the ECAL volume can improve signal efficiency, though cost considerations must also be addressed.
@@
Figure 6:  The distribution of recoil angles for different processes. Recoil angle is defined as the angle between the particle’s momentum direction and the beam direction. In this figure, the ECAL has a cross-sectional area of 52.5$\times$52.5 cm2. Dark photon signals exhibit larger recoil angles compared to inclusive background processes. As the dark photon mass increases, more events tend to impact the peripheral regions of the ECAL.
@@
A signal region for the calorimeter is defined in Table 1: the total energy in the ECAL is less than 2.5 GeV, the total energy in the HCAL is less than 30 MeV, and the maximum energy of a single unit in the HCAL is less than 0.1 MeV. This signal region is designed to achieve low background within $3\times 10^{14}$ electrons-on-target events. It is derived from the combined analysis of all sub-detectors, using thresholds for the ECAL and other sub-detectors, and applying extrapolation methods to exclude all backgroundsChen:2022liu ; Wang2024 . The optimization results of the ECAL are evaluated by signal efficiency, which is defined as the ratio of the number of events entering the signal region to the number of total simulated events.
Table 1:  Signal region for calorimeters.
@@
Figure 7: Signal efficiency as a function of dark photon mass for various ECAL configurations. The signal efficiency is defined as the ratio of the number of events entering the signal region in Table 1, to the number of total simulated events. In the legends, the format "x×$\times$×y×$\times$×z" represents the number of crystals in the ECAL along the transverse-x, transverse-y, and longitudinal-z dimensions, with each crystal measuring 2.5$\times$2.5$\times$4.0 cm3. The beam travels along the longitudinal direction. (a) Optimization of the transverse size of the ECAL, showing signal efficiency for different transverse dimensions. Increasing the ECAL’s transverse size can enhance signal efficiency, particularly for higher dark photon masses. (b) Optimization of the longitudinal size of the ECAL, showing signal efficiency for different longitudinal dimensions. It shows that increasing the ECAL’s longitudinal size improves signal efficiency.
The ECAL size was optimized in two dimensions: transverse dimension and longitudinal dimension, with the longitudinal direction being aligned along the beam direction. During optimization, the individual crystal size remains constant at 2.5$\times$2.5$\times$4.0 cm3, which is determined by the maximum density that our electronics can handle guo2024 .
First, the ECAL’s transverse dimensions were adjusted while keeping the longitudinal length constant at 11 layers. Four different calorimeter sizes were simulated. Results in Figure 7(a) show that as the transverse size increases, signal efficiency for different dark photon masses also increases. The average signal efficiency for all mass points under each size is listed in Table 2.
@@
V Energy distribution
As a fixed target experiment, incident particles without significant energy loss primarily hit the central area of the ECAL, resulting in substantial energy deposition in the central region’s crystals in the first few layers. Additionally, the energy absorbed by crystals in different regions varies greatly, with those closer to the edges typically experiencing minimal energy deposition. To determine the energy dynamic range for the future detector and explore potential triggering methods for the ECAL, the energy deposition in crystals from different regions was investigated.
V.1 Energy deposition in different regions
With the optimized ECAL volume described in IV, the energy deposition in crystals for dark photon signals and background processes is shown in Figure 8. Inclusive background processes typically lead to larger energy deposits in the crystals, reaching up to 4 GeV. While the energy deposition from dark photon signals is slightly lower, with smaller mass dark photons depositing more energy. A dark photon with a mass of 1 MeV can deposit up to 3.5 GeV in the crystals, as dark photons have a very low probability of carrying minimal kinetic energy, causing the recoiling electron to retain most of the energy.
Figure 8:  The energy deposited in individual crystals by an 8 GeV electron beam incident on the ECAL. For inclusive background processes, the energy absorbed by a single crystal can reach up to 4 GeV, while for dark photon signal, it can reach up to 3.5 GeV.
Figure 9:  ECAL divided into four regions, from the center to the periphery. Red: region-I, green: region-II, blue: region-III, yellow: region-IV. The beam enters the center of the detector along the direction from left to right in the diagram. The beam travels from left to right in the figure, with the beam direction defined as the longitudinal direction (z).
@@ V.1 Energy deposition in different regions
The ECAL was divided into four sections (Figure 9) from the center to the periphery to study energy deposition in different regions. The number of crystals in each region is shown in Table 4. Region-I (red) is the core area of shower development, containing 125 crystals each with a volume of 2.5$\times$2.5$\times$4 cm3, approximately equivalent to 18 radiation lengths. It covers the entire beam spot with a radius of 3 cm and an additional area of one Molière radius. The energy absorbed by each crystal in these four regions is shown in Figure 10. Crystals in the central and near-central regions absorb significantly more energy. In contrast, crystals in Region-III and Region-IV, farther from the shower center, absorb noticeably less energy. For dark photon signals, the energy deposition is even lower than for inclusive background. Figure 11 shows that crystals in Region-I, with only 2.6% of the ECAL volume, absorb more than 90% of the energy for inclusive backgrounds.
@@
Figure 10:  Energy absorbed by single crystals in the four regions shown in Figure 9. Crystals in the central regions (Region-I and Region-II) absorb significantly more energy, while the crystals in Region-III and Region-IV, farther from the shower center, absorb noticeably less energy.
Figure 11:  Energy ratio in different regions shown in Figure 9 for various processes. Energy ratio is defined as the ratio of the total energy deposited in the crystals within a specific region ($E_{i}^{Total}$) to the total energy deposited across all crystals in the ECAL ($E_{ECAL}^{Total}$). Both signal and background processes deposit the most energy in Region-I. However, for dark photon processes, as the dark photon mass increases, a larger proportion of the energy is absorbed in the outer regions.
V.2 Energy limits on channels
The setup of the energy dynamic range directly impacts the accuracy of energy measurements. Generally, a smaller dynamic range allows for higher measurement precision. To identify the minimum dynamic range that satisfies the energy measurement requirements of the ECAL, an energy limit was imposed on each crystal. When the energy deposited in a crystal exceeds this limit, it is capped at the limit, mimicking the saturation behavior observed in a real detector. However, if the energy limit is set too low, it can reduce the total energy measured by the ECAL, potentially causing background events with substantial energy loss to be misidentified as dark photon signals. Since our focus is on events characterized by significant energy loss, we investigated how the energy limit influences the number of events where the total energy measured by the ECAL falls below 4 GeV.
Figure 12:  ECAL energy distribution of inclusive background after applying different energy limits to channels. The energy limit represents the maximum energy that a single channel can measure, analogous to the saturation in detector electronics. When the limit is set too low, the energy measured by the ECAL can deviate significantly from the incident energy, shifting to a lower energy region. This shift can lead to the misidentification of background events as dark photon signals.
Figure 12 shows the total energy distribution in the ECAL after applying different energy limits to channels. As the energy limit decreases to below 3 GeV, the total energy measured by the ECAL starts to decrease. This is comparable to the saturation effect caused by the insufficient dynamic range of the electronics in detector. If the energy limit is greater than 3 GeV, the energy measured by the ECAL is almost unaffected by the limit. To ensure the accuracy of the ECAL’s energy measurement, the dynamic range of individual channels should not impact the total energy. From Figure 12, a dynamic range greater than 3 GeV is appropriate.
Table 5:  Ratios between the number of events with ECAL energy less than 4 GeV and the number of total simulated events, calculated after applying various channel energy limits to background and dark photon signals of different masses. As the limit decreases, the energy measured by the ECAL may also decrease, leading to more events falling into the region where the total ECAL energy is less than 4 GeV.
@@ V.2 Energy limits on channels
Table 5 shows the ratios between the number of events with ECAL energy less than 4 GeV and the number of total simulated events. These ratios were derived from background and dark photon signals of varying masses after applying different channel energy limits. As the energy limit decreases, the inclusive background has increasing energy loss, resulting in a growing number of cases where the ECAL energy is less than 4 GeV. The results indicate that a dynamic ranges of 100 MeV or 500 MeV is insufficient, as they demonstrate poor background rejection capability with a statistical sample of one million events. When the energy limit on channels is set to above 500 MeV, nearly all events deposit more than 4 GeV of energy in the ECAL, except for two photon-muon pair processes and one nuclear process, which can be effectively vetoed by the HCAL. This indicates very few crystals exhibit energy depositions exceeding 1 GeV, which is also reflected in the normalized entries shown in Figure 8. Therefore, from the perspective of background rejection, the energy limit should be set to at least 1 GeV.
Overall, a 4 GeV energy dynamic range is optimal for DarkSHINE ECAL, as it does not impact the accuracy required for energy measurements while still ensuring a great capability of background rejection. The readout electronics in future detectors can be designed based on this standard.
V.3 Trigger efficiency
In the high-frequency beam environment of DarkSHINE, the majority of events collected by the detector are background events that do not interest us. These background events can significantly strain data transmission and storage on the backend. However, by implementing an online trigger to directly filter out events that clearly don’t match the characteristics of a dark photon signals, we can save a considerable amount of resources. A preliminary design for such a trigger involves summing the energy across selected channels in ECAL. If the summed energy is less than 4 GeV, the event will be saved. A trigger efficiency is defined as the ratio of the number of triggered events, where the summed energy in the trigger region is below 4 GeV, to the total number of events. Table 6 shows the trigger efficiency across different trigger regions.
Table 6:  Table of trigger efficiency across different trigger regions. The trigger efficiency is defined as the ratio of the number of triggered events, where the summed energy in the trigger region is below 4 GeV, to the total number of events. The trigger regions I, II, III and IV correspond to those shown in Figure 9.